# Topological classification of black holes surrounded by quintessence

### v1

@@ Abstract
In this paper, we consider a class of black hole solutions surrounded by quintessence as defects possessing topological characteristics within the thermodynamic space. By formulating a generalized off-shell free energy and calculating the winding number for each black hole branch through the residues of isolated first-order poles of characteristic function defined by this free energy, the winding number reflects the local properties of black holes. The topological number, being the sum of winding numbers for all black hole branches, reflects global properties. Utilizing the residue method, we discovered that the topological numbers for Schwarzschild, Kerr, and Kerr-AdS black holes surrounded by quintessence are consistent with those not surrounded by quintessence. Furthermore, 4D/5D Einstein-Gauss Bonnet black holes under a quintessence field belong to the same topological class as Kerr black holes. Thus both the rotation parameter and Gauss-Bonnet parameter in this background field affect the topological number of the black hole. Importantly, the winding numbers for generation and annihilation points can be calculated using the residues of second-order poles of the characteristic function.
@@ I Introduction
Black hole thermodynamics, as a crucial avenue for exploring the deep connections between black hole properties and thermodynamic principles, originates from a groundbreaking realization: black holes possess thermodynamic quantities such as entropy and temperature  Hawking1 ; Hawking2 ; Hawking3 ; Hawking4 . Although substantial progress has been made in this field, a unified and systematic framework for characterizing the intrinsic features of various black hole systems is still lacking. In recent years, the emergence of the thermodynamic topology approach has provided a novel perspective. This method treats black holes as topological defects in thermodynamic space and defines topological invariants by analyzing the singularity structure of generalized free energy. Based on this framework, most black holes can be classified into three distinct topological classes: +1, 0, and –1, which reflect their global stability Wei2 . Due to its generality and computational efficiency, this approach has been widely applied to black holes in diverse settings, including rotating AdS black holes T4 ; T5 , phantom AdS black holes chen , Lovelock AdS black holes  T6 ; T7 , Born-Infeld black holes  T9 , as well as more complex configurations such as nonlinearly charged black holes  T10 , R-charged black holes T11 and hairy black holes  T8 . Furthermore, topology has also been extended to the study of black holes in Rényi statistics  T12 ; T13 . This topological framework opens new avenues for understanding the geometric nature of gravity. Naturally, it has been extended to black holes in the presence of dark energy fields.
Dark Matter (DM) and Dark Energy (DE) are two prominent issues in modern cosmology and astrophysics. Investigating these components facilitates understanding of the structure and evolution of the universe. Astronomical observations indicate that the universe is expanding at an accelerated rate due to DE, which involves a negative pressure driving the expansion. The introduction of DE has provided a reasonable explanation for this accelerated expansion Q1 ; Q2 ; Q3 . Statistically, DE accounts for about 73% of the total energy content of the universe, DM for about 23%, and ordinary baryonic matter only about 4%. Despite extensive astronomical observations, the fundamental nature and origin of DE remain a mystery. Various models have been proposed to explain DE, such as quintessence and the cosmological constant. Quintessence is characterized by an equation relating pressure $P$ and energy density $\rho_{q}$, expressed as $P=\omega_{q}\rho_{q}$. Depending on the value of the quintessence state parameter $\omega_{q}$, three different scenarios arise: phantom energy ($\omega_{q}<-1$), the cosmological constant ($\omega_{q}=-1$), and quintessence-like matter ($-1<\omega_{q}<-1/3$).
@@
This work investigates the thermodynamic properties of typical black holes in such a dark energy field from a topological perspective. The structure of the paper is as follows: Section II introduces the topological number of Schwarzschild black holes with a quintessence background; Section III analyzes Kerr and Kerr-AdS black holes in this context; Section IV discusses the effect of the quintessence field on the thermodynamic topological number of EGB black holes; the final section provides our conclusions.
II Schwarzschild black hole
@@
$N(r)=1-\frac{2M}{r}-\frac{c}{r^{3\omega_{q}+1}},$ (2)
here, $M$ denotes the black hole mass, and $c$ is a positive normalization factor linked to the quintessence energy density. When the equation-of-state parameter lies within the interval $-1<\omega_{q}<-1/3$, quintessence effectively accounts for the observed cosmic acceleration. Note that when $c=0$ the metric reduces to the standard Schwarzschild black hole metric. In this section, we focus on a typical dark energy model and adopt the representative parameter $\omega_{q}=-2/3$. For this quintessence model, the energy-momentum tensor takes the following form:
@@
Next, we use the residue theorem to study the topological number of the Schwarzschild black hole surrounded by a quintessence. according to eqs. (5-7), the form of the generalized Helmholtz free energy is written as Wei2
$\mathcal{F}$ $=M-\frac{S}{\tau}$ (8)
$=\frac{1}{2}(r_{+}-cr_{+}^{-3\omega_{q}})-\frac{\pi r_{+}^{2}}{\tau},$
$\tau$ denotes an auxiliary parameter that can be interpreted as the inverse temperature parameter of the cavity surrounding the black hole. The free energy corresponds to its on-shell value when $\tau=1/T$. According to the $\phi$-mapping theory Duan , a vector field is defined as
$\phi=(\partial_{r_{+}}\mathcal{F},-\cot\Theta\csc\Theta),$ (9)
$\Theta$ is an auxiliary factor introduced for convenience in topological calculations. The value $\Theta=\pi/2$ orresponds to the zero point of the vector field $\phi$. The components of the vector field $\phi$ can be written as follows:
@@
$\tau=\frac{4\pi r_{+}^{2+3\omega_{q}}}{r_{+}^{1+3\omega_{q}}+3c\omega_{q}}=\mathcal{G}(z).$ (12)
where $\mathcal{G}(z)$ denotes a function of the complex variable $z$. Subsequently, we construct the rational complex function $\mathcal{R}(z)$ as follows:
$\mathcal{R}(z)=\frac{1}{\tau-\mathcal{G}(z)},$ (13)
according to eq. (12) and eq. (13), we can obtian
$\mathcal{R}(z)=-\frac{z^{1+3\omega_{q}}+3c\omega_{q}}{4\pi z^{2+3\omega_{q}}-\tau(z^{1+3\omega_{q}}+3c\omega_{q})}.$ (14)
Considering the typical parameter $\omega_{q}=-2/3$, it is thus derived
$\mathcal{R}(z)=\frac{1-2cz}{(1-2cz)\tau-4\pi z}=\frac{1-2cz}{\mathcal{A}(z)}.$ (15)
The analysis of $\mathcal{A}(z)=0$ shows that there is only one positive real root, $z_{1}=\frac{\tau}{2(2\pi+c\tau)}$, According to the residue theorem, $z_{1}$ is a simple isolated pole of the function $\mathcal{R}(z)$, and the residue of $\mathcal{R}(z)$ is given by $\operatorname{Res}\mathcal{R}(z_{1})=\lim_{z\to z_{1}}(z-z_{1})\mathcal{R}(z)$. Thus, we can compute the residue of the characteristic complex function $\mathcal{R}(z)$ at the singularity $z_{1}$ as
$\operatorname{Res}\mathcal{R}(z_{1})=-\frac{1-2cz_{1}}{2(2\pi+c\tau)},$ (16)
It is worth noting that the sign of the numerator in the above expression plays an important role in our analysis, as it directly affects the sign of the winding number. According to the condition for an extremal black hole, it is straightforward to determine that $01-2cr_{+}>$ 0. Based on the definition of the winding number equation Q21
$w_{i}=\frac{\operatorname{Res}\mathcal{R}\left(z_{i}\right)}{\left|\operatorname{Res}\mathcal{R}\left(z_{i}\right)\right|}=\operatorname{Sgn}\left[\operatorname{Res}\mathcal{R}\left(z_{i}\right)\right],$ (17)
Thus, the winding number of $z_{1}$ and the total topological number of the system are given by
$w_{1}=-1,W=-1.$ (18)
By employing the framework of topological current theory, Figure 1 presents the zero point of the component $\phi^{r_{+}}$ and the corresponding unit vector field $n$. Here $r_{0}$ denotes an arbitrary length scale determined by the size of the cavity enclosing the black hole. The figure shows that the Schwarzschild black hole surrounded by quintessence possesses a single unstable branch, characterized by a topological charge of $W=-1$. This result suggests that the winding number and topological number of the Schwarzschild black hole remain unchanged despite the influence of quintessence, aligning with findings for the standard Schwarzschild case Wei2 . In other words, the quintessence field does not modify the black hole’s topological classification. Moreover, we find that the topological number obtained via the residue theorem is consistent with that derived from the topological current method. In the following, we will further examine this conclusion by presenting additional examples.
(a) The zero points of $\phi^{r_{+}}$ in the $\tau-r_{+}$ plane
(b) The unit vector field $n$ on the $r_{+}-\Theta$ plane
Figure 1: The topological properties of the Schwarzschild black hole surrounded by quintessence, we set $cr_{0}^{2}=0.1$.
III Kerr black hole and it’s AdS case
This section investigates the topological classification of Kerr spacetime under a quintessence field, along with its extension to an AdS background. First, in the Boyer-Lindquist coordinate system, the Kerr spacetime metric in the quintessence background is given by: kbh
@@ III Kerr black hole and it’s AdS case
Based on these thermodynamic quantities, the expression for the free energy can be derived
@@
$\phi^{r_{h}}$ $=\frac{1}{2}(1-\frac{a^{2}}{r_{+}^{2}}+3c\omega_{q}r_{+}^{-1-3\omega_{q}})-\frac{2\pi r_{+}}{\tau},$ (25)
@@
$\tau=\frac{4\pi r_{+}^{3+3\omega_{q}}}{r_{+}^{2+3\omega_{q}}-a^{2}r_{+}^{3\omega_{q}}+3c\omega_{q}r_{+}}.$ (26)
Similar to the previous section, the characteristic function in this case is defined as
$\mathcal{R}(z)=\frac{z^{2+3\omega_{q}}-a^{2}z^{3\omega_{q}}+3cz\omega_{q}}{4\pi z^{3+3\omega_{q}}-\tau(z^{2+3\omega_{q}}-a^{2}z^{3\omega_{q}}+3cz\omega_{q})}.$ (27)
When $\omega_{q}=-2/3$, we simplify the above equation to
$\mathcal{R}(z)=-\frac{z^{2}-2cz^{3}-a^{2}}{4\pi z^{3}-\left(z^{2}-2cz^{3}-a^{2}\right)\tau}=-\frac{z^{2}-2cz^{3}-a^{2}}{\mathcal{A}(z)}.$ (28)
$\mathcal{A}(z)$ is regarded as a polynomial function, applying the fundamental theorem of algebra, that is, any polynomial
$y(x)=m_{0}+m_{1}x+\cdots+m_{n}x^{n},\quad m_{n}\neq 0,$ (29)
can be decomposed into
$y(x)=m_{n}\left(x-z_{1}\right)\left(x-z_{2}\right)\cdots\left(x-z_{n}\right),$ (30)
where $z_{i}$ is the root of $y(x)$,which can be a complex number or a real number. Therefore, the roots of $\mathcal{A}(z)$ correspond to the poles of the function $\mathcal{R}(z)$ and the equation  (28) can be rewritten as
$\mathcal{R}(z)=-\frac{z^{2}-2cz^{3}-a^{2}}{(z-z_{1})(z-z_{2})(z-z_{3})\dots(z-z_{n})}.$ (31)
(a) The zero points of $\phi^{r_{+}}$ in the $\tau-r_{+}$ plane, a creation point located in $\tau/r_{0}=\tau_{c}/r_{0}=3.4438$.
(b) The roots of $\mathcal{A}(z)$.
(c) The unit vector field $n$ of $\tau/r_{0}=5$.
Figure 2: The topological properties of the kerr black hole surrounded by quintessence, we set $a/r_{0}=0.1$, $cr_{0}^{2}=0.1$.
@@
The plot of $r_{+}$ versus $\tau$ is shown in Fig. 2(a), where a stable small black hole branch and an unstable large black hole branch can be observed. At the point $\tau_{c}=\frac{12\sqrt{3}\pi a}{2-6\sqrt{3}ca}$, we find that $(d^{2}\tau/dr_{+}^{2})>0$, indicating that $\tau_{c}$ is a generation point. From Fig. 2(b), it can be seen that when $\tau>\tau_{c}$, the polynomial has one negative real root and two positive real roots

### v2

# Universal topological classes of black holes surrounded by quintessence

@@ Abstract
This work explores the universal topological classes of various black hole configurations immersed in a quintessence field. The results indicate that the Schwarzschild black hole surrounded by quintessence is of the $W^{1-}$ type, exhibiting thermodynamic instability in both extremal temperature limits. The introduction of rotation alters its topological nature to the $W^{0+}$ class, where the Kerr black hole displays coexistence of a stable small black hole and an unstable large black hole at low temperatures. In comparison, anti de Sitter (adS) black holes display different thermodynamic topologies. The Schwarzschild–adS solution corresponds to the $W^{0-}$ class, where an unstable small branch and a stable large branch emerge in the high temperatures. Meanwhile, the Kerr–adS classified as $W^{1+}$, maintains stability in both high- and low-temperature limits. Furthermore, the 4D/5D quintessential black hole belongs to the $W^{0+}$ class within the Einstein–Gauss–Bonnet framework. Overall, parameters such as rotation and Gauss-Bonnet in quintessence background field affect the topological classifications of the black holes, while the quintessence field primarily modifies the stability range without changing the topological class itself.
@@ I Introduction
Black hole thermodynamics, as a crucial avenue for exploring the deep connections between black hole properties and thermodynamic principles, originates from a groundbreaking realization: black holes possess thermodynamic quantities such as entropy and temperature H1 ; H2 ; H3 ; H4 . Although substantial progress has been made in this field, a unified and systematic framework for characterizing the intrinsic features of various black hole systems is still lacking. In recent years, the emergence of the thermodynamic topology approach has provided a novel perspective. This method treats black holes as topological defects in thermodynamic space and defines topological invariants by analyzing the singularity structure of generalized free energy. Based on this framework, most black holes can be classified into three distinct topological classes: +1, 0, and –1, which reflect their global stability Wei2 . Due to its generality and computational efficiency, this approach has been widely applied to black holes in diverse settings, including rotating AdS black holes T4 ; T5 , phantom AdS black holes chen , Lovelock AdS black holes T6 ; T7 , Born-Infeld black holes T9 , as well as more complex configurations such as nonlinearly charged black holes T10 , R-charged black holes T11 and hairy black holes T8 . Furthermore, topology has also been extended to the study of black holes in Rényi statistics T12 ; T13 . Recently, Ref. T14 further refined the topological classification of black hole thermodynamics into four universal topological types. Subsequently, Ref. T15 expanded the existing classification framework. The new topological classification method has already been applied to other types of black holes T16 ; T17 ; T18 ; T19 . This topological framework opens new avenues for understanding the geometric nature of gravity. Naturally, it has been extended to black holes in the presence of dark energy fields.
Dark Matter (DM) and Dark Energy (DE) are two prominent issues in modern cosmology and astrophysics. Investigating these components facilitates understanding of the structure and evolution of the universe. Astronomical observations indicate that the universe is expanding at an accelerated rate due to DE, which involves a negative pressure driving the expansion. The introduction of DE has provided a reasonable explanation for this accelerated expansion Q1 ; Q2 ; Q3 . Statistically, DE accounts for about 73% of the total energy content of the universe, DM for about 23%, and ordinary baryonic matter only about 4%. Despite extensive astronomical observations, the fundamental nature and origin of DE remain a mystery. Various models have been proposed to explain DE, such as quintessence and the cosmological constant. Quintessence is characterized by an equation relating pressure $P$ and energy density $\rho_{q}$, expressed as $P=\omega_{q}\rho_{q}$. Depending on the value of the quintessence state parameter $\omega_{q}$, three different scenarios arise: phantom energy ($\omega_{q}<-1$), the cosmological constant ($\omega_{q}=-1$), and quintessence-like matter ($1<\omega_{q}<-1/3$).
@@
The structure of this paper is organized as follows: in the next section, we investigate the topological class of Schwarzschild black hole surrounded by quintessence. Section III shifts the focus to the thermodynamic topological classification of Kerr black hole in the presence of quintessence. Sections IV and V examine the Schwarzschild–AdS and Kerr–AdS black holes within the quintessence matter, respectively. Section VI discusses the topological characteristics of Einstein–Gauss–Bonnet (EGB) black holes. The final section provides our conclusions.
II Schwarzschild black hole surrounded by quintessence
@@
$f(r)=1-\frac{2M}{r}-\frac{c}{r^{3\omega_{q}+1}},$ (2)
here, $M$ denotes the black hole mass, and $c$ is a positive normalization factor linked to the quintessence energy density. When the equation-of-state parameter lies within the interval $1<\omega_{q}<-1/3$, quintessence effectively accounts for the observed cosmic acceleration. Note that when $c=0$ the metric reduces to the standard Schwarzschild black hole metric. In this paper, we focus on a typical dark energy model and adopt the representative parameter $\omega_{q}=-2/3$. For this quintessence model, the energy-momentum tensor takes the following form
@@
It is straightforward to confirm that, for the Schwarzschild black hole surrounded by quintessence (Sch–BH-$\omega_{q}$), the Hawking temperature tends to vanish as the event horizon radius $r_{+}$ approaches infinity, whereas it becomes divergent when $r_{+}$ approaches the minimum horizon radius $r_{m}$. As a result, the inverse temperature $\beta(r_{+})$ shows
$\beta(r_{m})=0,\beta(r_{\infty})=\infty.$ (8)
At above asymptotic limits, defect curve $\beta(r_{+})$ is analytical in the range $(r_{m},\infty)$. The topological properties of the Sch–BH-$\omega_{q}$ are fundamentally rooted in the concept of generalized off shell free energy. According to eqs. (5-7), the form of the free energy is written as Wei2
$\mathcal{F}$ $=M-\frac{S}{\tau}=\frac{1}{2}(r_{+}-cr_{+}^{-3\omega_{q}})-\frac{\pi r_{+}^{2}}{\tau},$ (9)
$\tau$ denotes an auxiliary parameter that can be interpreted as the inverse temperature parameter of the cavity surrounding the black hole. The free energy corresponds to its on-shell value when $\tau=\beta=1/T$. According to Duan’s $\phi$-mapping theory Duan1 ; Duan2 , we introduce an additional parameter $\Theta\in(0,\pi)$, accordingly, the two-component vector $\phi$ is specified by the gradient of as $\tilde{\mathcal{F}}=\mathcal{F}+\frac{1}{\sin\Theta}$ T14
[…]
$\Theta$ is an auxiliary factor introduced for convenience in topological calculations. The value $\Theta=\pi/2$ orresponds to the zero point of the vector field $\phi$. Based on $\phi$-mapping theory, each black hole state is assigned a winding number reflecting its topological charge. Locally stable states correspond to $w=+1$, while unstable ones have w=1w=-1. The sum of all local winding numbers defines the global topological number $W$, which determines the black hole’s overall topological class.
We establish the components of the vector field $\phi$ as:
@@
$\tau=\beta=\frac{4\pi r_{+}^{2+3\omega_{q}}}{r_{+}^{1+3\omega_{q}}+3c\omega_{q}}.$ (13)
Figure 1: The zero points of $\phi^{r_{+}}$ in the $r_{+}-\beta$ plane for Sch-BH-$\omega_{q}$, we set $a/r_{0}=0.1$, $cr_{0}^{2}=0.1$.
Figure 1 presents the component $\phi^{r_{+}}$ in the $r_{+}-\beta$ plane, here $r_{0}$ denotes an arbitrary length scale determined by the size of the cavity enclosing the black hole. The figure shows that the Sch–BH-$\omega_{q}$ possesses a single unstable branch. We proceed to examine the variation of the vector field $\phi$ (see Fig. 2) along the boundary defined by Eq. (8). The complete boundary is outlined by the closed contour $C=I_{1}\cup I_{2}\cup I_{3}\cup I_{4}$, whose segments and overall shape are shown in Fig. 3, encompassing the entire parameter domain. Due to the construction of $\phi$, it remains perpendicular to $I_{2}$ and $I_{4}$ T14 , implying that the key asymptotic properties are determined by its behavior on $I_{1}$ and $I_{3}$. Specifically, as $r_{+}\to r_{m}$, the vector field directs rightward with an inclination governed by $\phi^{\Theta}$; whereas in the limit $r_{+}\to\infty$, it turns leftward, as illustrated in Fig. 2 and Fig. 3. Table I presents the orientations of the vector field $\phi$ along segments $I_{1}$ and $I_{3}$.
[…]
Figure 2: The unit vector field $n$ on the $r_{+}-\Theta$ plane for the Sch-BH-$\omega_{q}$, we set $\beta/r_{0}=20$.
Figure 3: The asymptotic behavior of the $n$-vector field $\phi$ at the boundary ($C=I_{1}\cup I_{2}\cup I_{3}\cup I_{4}$).
Figure 4: Contour $\Phi_{1}$ depicts the variation of the vector field $\phi$ components along the pat $C_{1}$ show in Fig. 2 for the Sch-BH-$\omega_{q}$ case.
Table 1: The orientation is indicated by the arrow of $\phi^{r_{+}}$ for the Sch-BH-$\omega_{q}$ and Kerr-BH-$\omega_{q}$ and the topological numbers.
BH spacetime $I_{1}$ $I_{2}$ $I_{3}$ $I_{4}$ $W$
Sch-BH-$\omega_{q}$ $\leftarrow$ $\uparrow$ $\to$ $\downarrow$ -1
Kerr-BH-$\omega_{q}$ $\leftarrow$ $\uparrow$ $\leftarrow$ $\downarrow$ 0
[…]
III Kerr black hole surrounded by quintessence
This section investigates the topological classification of Kerr spacetime under a quintessence field. First, in the Boyer-Lindquist coordinate system, the Kerr spacetime metric in the quintessence background is given by: kbh
@@ III Kerr black hole surrounded by quintessence
the asymptotic behavior of the inverse temperature $\beta(r_{+})$ for the Kerr black hole surrounded by quintessence (Kerr-BH-$\omega_{q}$) has the limits
$\beta(r_{m})=\infty,\beta(r_{\infty})=\infty.$ (19)
Based on above thermodynamic quantities, the expression for the free energy can be derived
@@
$\phi^{r_{+}}$ $=\frac{1}{2}(1-\frac{a^{2}}{r_{+}^{2}}+3c\omega_{q}r_{+}^{-1-3\omega_{q}})-\frac{2\pi r_{+}}{\tau},$ (21)
@@
$\tau=\beta=\frac{4\pi r_{+}^{3+3\omega_{q}}}{r_{+}^{2+3\omega_{q}}-a^{2}r_{+}^{3\omega_{q}}+3c\omega_{q}r_{+}}.$ (22)
Figure 5: The zero points of $\phi^{r_{+}}$ in the $r_{+}-\beta$ plane for Kerr-BH-$\omega_{q}$, we set $a/r_{0}=0.1$, $cr_{0}^{2}=0.1$.
Figure 6: The unit vector field $n$ on the $r_{+}-\Theta$ plane for the Kerr-BH-$\omega_{q}$ for $\beta/r_{0}=5$.
Figure 7: Contours $\Phi_{2}$ and $\Phi_{3}$ depict the variation of the vector field $\phi$ components along the pats $C_{2}$ and $C_{3}$ show in Fig. 6 for the Kerr-BH-$\omega_{q}$ case.
@@
The plot of $r_{+}$ versus $\beta$ is shown in Fig. 5, where a small black hole branch and a large black hole branch can be observed. A special point $\beta_{c}=\frac{12\sqrt{3}\pi a}{2-6\sqrt{3}ca}$ for which $d\beta/dr_{+}=0$ is degenerate, at this point we find that $(d^{2}\beta/dr_{+}^{2})>0$, indicating that $\beta_{a}$ is a generation point. When we set $a/r_{0}=0.1$, $cr_{0}^{2}=0.1$, the generation point located in $\beta/r_{0}=\beta_{c}/r_{0}=3.4438$.
The characteristics of the vector field $n$ shown in Fig. 6, together with its asymptotic behavior along the boundary segments $I_{i}$ (analogous to the case of the Sch-BH-$\omega_{q}$ illustrated in Fig. 3), are summarized in Table I. It is found that in the limits $r_{h}\to r_{m}$ or $r_{h}\to\infty$, the vector field plot of $\phi$ points leftward.
Figure 7 illustrates the variation pattern of the vector field $\phi$ along the chosen contours. It can be observed that the closed trajectory $\Phi_{2}$ rotates counterclockwise, leading to winding numbers of +1 and -1 for the first and second zero points, respectively. This result indcates a total topological number $W=0$ for the Kerr-BH-$\omega_{q}$.
For the universal thermodynamic behavior, when $\beta\to\infty$, both a stable small black hole and an unstable large black hole coexist. In contrast, as $\beta\to 0$, the black hole phase disappears entirely. Therefore, the Kerr-BH-$\omega_{q}$ is identified with the topological class $W^{0+}$. Moreover, the winding numbers and universal topological class for the Kerr-BH-$\omega_{q}$ match those of the standard Kerr case T17 , yet they deviate from the values of the Schwarzschild black hole under the same dark energy influence. Therefore, the spacetime rotation in the quintessence framework influences its topological characteristics.
IV Schwarzschild-adS black hole surrounded by quintessence
Then using the ideas of Kiselev who proposed the solution of the Schwarzschild–anti de Sitter black hole surrounded by quintessence (Sch-aBH-$\omega_{q}$) Q4
$ds^{2}=-f(r)dt^{2}+\frac{dr^{2}}{f(r)}+r^{2}d\theta^{2}+r^{2}\sin^{2}\theta d\phi^{2},$ (23)
with
$f(r)=1-\frac{2M}{r}-\frac{\Lambda r^{2}}{3}-\frac{c}{r^{3\omega_{q}+1}},$ (24)
where $\Lambda$ is the cosmological constant. Then by using Eq. (24), one gets
$M=\frac{r_{+}}{2}-\frac{cr_{+}^{-3\omega_{q}}}{2}-\frac{\Lambda r_{+}^{3}}{6}.$ (25)
The Hawking temperature and entropy associated with the black hole horizon can be determined
$T=\frac{1}{4\pi r_{+}}-\frac{\Lambda r_{+}}{4\pi}+\frac{3c\omega_{q}r_{+}^{-2-3\omega_{q}}}{4\pi},$ (26)
$S=\pi r_{+}^{2}.$
The asymptotic behavior of the inverse temperature $\beta(r_{+})$ exhibits
$\beta(r_{m})=0,\penalty 10000\ \beta(r_{\infty})=0.$ (27)
The generalized off-shell Helmholtz free energy can be easily obtained
$\mathcal{F}=\frac{r_{+}}{2}-\frac{cr_{+}^{-3\omega_{q}}}{2}-\frac{\Lambda r_{+}^{3}}{6}-\frac{\pi r_{+}^{2}}{\tau},$ (28)
the components of the vector $\phi$ can be computed as
$\phi^{r_{+}}=\frac{1}{2}+4\pi Pr_{+}^{2}+\frac{3c\omega_{q}r_{+}^{-1-3\omega_{q}}}{2}-\frac{2\pi r_{+}}{\tau},$ (29)
$\phi^{\theta}=-\cot\Theta\csc\Theta,$ (30)
and the inverse temperature is
$\beta=\frac{4\pi r_{+}^{2+3\omega_{q}}}{r_{+}^{1+3\omega_{q}}+8\pi Pr_{+}^{3+3\omega_{q}}+3c\omega_{q}}.$ (31)
Figure 8: The zero points of $\phi^{r_{+}}$ in the $r_{+}-\beta$ plane for the Sch-aBH-$\omega_{q}$, we set $a/r_{0}=0.1$, $cr_{0}^{2}=0.1$, and $Pr_{0}^{2}=0.1$.
Figure 9: The unit vector field $n$ on the $r_{+}-\Theta$ plane for the Sch-aBH-$\omega_{q}$, we set $\beta/r_{0}=3$.
Figure 10: Contours $\Phi_{4}$ and $\Phi_{5}$ depict the variation of the vector field $\phi$ components along the pats $C_{4}$ and $C_{5}$ show in Fig. 9 for the Sch-aBH-$\omega_{q}$ case.
In this framework, the pressure term is defined by $P=-\Lambda/8\pi$. The plot of $r_{+}$ versus $\beta$ is shown in Fig. 8. At the point $\beta_{c}=\frac{2\pi}{3\sqrt{2\pi P}-c}$ which is a degenerate point, we find that $(d^{2}\beta/dr_{+}^{2})<0$, indicating that $\beta_{c}$ is an annihilation point. When $\beta<\beta_{c}$, two black hole phases coexist a small branch and a large branch. Once $\beta>\beta_{c}$, these phases disappear, leaving no viable black hole solution. The annihilation point locates at $\beta_{c}/r_{0}=4.23016$ for $cr_{0}^{2}=0.1$ and $Pr_{0}^{2}=0.1$.
The corresponding vector field $n$ associated with the Sch-aBH-$\omega_{q}$ is illustrated in Fig. 9. As shown in the figure, this black hole exhibits two distinct zero points (depicted as black dots), which is in contrast to the Sch-BH-$\omega_{q}$ that possesses only one. The asymptotic orientation of the vector field $\phi$ at the boundaries $I_{i}$, as listed in Table II, reveals that as $r_{+}\to r_{m}$ or $r\to\infty$, the field tends to point rightward. This directional behavior indicates that the thermodynamic topology of the system corresponds to a particular universal class.
To further elucidate the topological structure, Fig. 10 presents the evolution of the vector field components ($\phi^{r_{+}},\phi^{\Theta}$) traced along the closed contours $C_{i}$ depicted in Fig. 9. The zero point of $\Phi_{i}$ is located at the origin. Contours with the first and second zeros possess winding numbers -1 and +1, respectively.